Q: solve for x and y :- (a+b)x+(a-b)y=2ab and (a+b)x-(a-b)y=ab .

Grade:10

2 Answers

Rinkoo Gupta
9 years ago
a+b)x+(a-b)y=ab
2ab(a+b)x-(a-b)y=ab
Adding both eqs, we get
(1+2ab)(a+b)x=2ab
x=2ab/(2ab+1)(a+b)
putting the value of x in eq (1)
(a+b).2ab/(2ab+1)(a+b) +(a-b)y=ab
(a-b)y=ab-2ab/(2ab+1)
=ab(2ab-1)/(2ab+1)
y=ab(2ab-1)/(2ab+1)(a-b)

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
kaushik kashyap
9 years ago
please check your answer in the second step , how do you get 2ab in the second step ?